# Predictive Inorganic Synthesis based on Machine Learning using Small Data sets: a case study of size-controlled Cu Nanoparticles

### v1

@@ Abstract
Copper nanoparticles (Cu NPs) have a broad applicability, yet their synthesis is sensitive to subtle changes in reaction parameters. This sensitivity, combined with the time- and resource-intensive nature of experimental optimization, poses a major challenge in achieving reproducible and size-controlled synthesis. While Machine Learning (ML) shows promise in materials research, its application is often limited by scarcity of large high-quality experimental data sets. This study explores ML to predict the size of Cu NPs from microwave-assisted polyol synthesis using a small data set of 25 in-house performed syntheses. Latin Hypercube Sampling is used to efficiently cover the parameter space while creating the experimental data set. Ensemble regression models, built with the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] successfully predict particle sizes with high accuracy (), outperforming classical statistical approaches (). Overall, this study highlights that, for lab-scale synthesis optimization, high-quality small datasets combined with classical, interpretable ML models outperform traditional statistical methods and are fully sufficient for quantitative synthesis prediction. This approach provides a sustainable and experimentally realistic pathway toward data-driven inorganic synthesis design.
@@ 1 Introduction
Metal nanoparticles (NPs) play a central role in a broad range of applications, including catalysis, electronics, and biomedicine, due to their unique size- and shape-dependent properties.[daniel:ChemInform2004, zhai:AdvancedFunctMat2010, gawande:ChemRev2016, pankhurst:JournalPhysD2003] However, controlling these properties requires very precise control of various NP characteristics, including their morphology (i.e., size and shape), composition, and surface chemistry down to the nanometer scale. Achieving this level of control remains one of the biggest challenges in NP synthesis, as small variations in the synthesis can lead to final particles with entirely different properties.[gawande:ChemRev2016, baig:MatAdv2021] Traditionally, synthesis optimization relies on iterative, empirical approaches that are slow, resource-intensive, and often lack mechanistic insight.[baig:MatAdv2021] To overcome these limitations, more systematic, data-driven strategies are needed, and recent advances in artificial intelligence (AI) and machine learning (ML) offer promising opportunities to accelerate the development of reproducible and tailored NP syntheses.[baig:MatAdv2021, baum:JournalChemInfMod2021, butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023, agrawal:APLMaterials2016]
@@
By uncovering hidden correlations in complex parameter spaces, ML can guide experimental design and reduce the number of required experiments.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] In materials science, ML is now routinely used to uncover the structure–property relationships of materials, thereby accelerating the discovery of new materials with specific properties.[butler:Nature2018, schmidt:NPJCompMat2019, sanchez-lengeling:Science2018, huang:Materials2023] However, for synthesis optimization, the key challenge lies in predicting how synthesis parameters determine the resulting crystal structure or particle morphology. These synthesis–structure relationships are far less explored, yet crucial for rational synthesis design. As a result, recent data-driven approaches are shifting towards models that learn synthesis–structure relationships to accelerate synthesis optimization.[agrawal:APLMaterials2016, vanpoucke:PolymerInt2022, williamson:InorgChem2023] Unlike traditional one-variable-at-a-time (OVAT) approaches, which are slow and fail to capture the multidimensional correlations that govern synthesis outcomes, ML can account for such interactions and thereby accelerate synthesis optimization.[williamson:InorgChem2023, lah:EngAnalysisBoundElements2024] However, while large data sets underpin many ML successes [Huo2022MLsynthesis, merchant-2023, curtarolo-2013], they are rarely feasible in laboratory-scale chemistry, where the creation of each data point requires a significant amount of time and resources. In the specific case of inorganic NP synthesis, literature-based data sets are often inconsistent and biased, as failed experiments remain unpublished[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024], as shown by Huo et al.[Huo2022MLsynthesis], who expected up to % of the values to be incorrect. As such, robust ML workflows for small, high-quality data sets often remain underexplored.
@@
ML can identify which synthesis parameters most strongly influence inorganic NP outcomes, as demonstrated by Guda et al.[guda:JournalPhysChemC2023], who trained ML models on 30 data points to classify the results of gold NP synthesis (success, shape, and formation kinetics) based on the precursor concentrations, reducing agent, and surfactant. ML also enables closed-loop experimentation, where ML models suggest subsequent experiments based on previous results, accelerating optimization and supporting high-throughput synthesis.[tao:AdvFunctMat2021, epps:AdvMat2020] Pellegrino et al.[pellegrino:ScientificRep2020] trained neural networks (NNs), with 20 initial data points, to predict the size, dispersion, and aspect ratio of NP, identifying optimal synthesis parameters through reverse engineering. Similarly, Williams et al.[williams:AppliedNano2023] combined Design of Experiments (DoE) and random forests to accurately predict the yield of Ag nanowires in a polyol-based flow reactor, based on 31 samples.
@@
In the current work, the prediction of copper (Cu) NP size as function of synthesis parameters is explored. Cu NPs hold significant promise for various catalytic processes, such as electro-catalysis and photo-catalysis, due to their abundance, cost-effectiveness, and versatile properties.[guo:AngewandteChemIE2014, reske:JACS2014, gawande:ChemRev2016] For example, Reske et al.[reske:JACS2014] showed that decreasing the size of Cu NPs significantly improves their catalytic activity. Different synthesis methods give rise to different types of Cu NPs, which may differ in particle size, morphology, and extent of oxidation. While metallic Cu NPs are nominally (), they often undergo (partial) surface oxidation depending on the synthesis conditions, leading to () phases of varying composition.[gawande:ChemRev2016] The wide variety of possible synthesis outcomes, as well as the large parameter space, complicates the development of robust synthesis procedures to obtain specific properties.
@@
Though ML is a tool with great potential for Cu NP synthesis, two challenges need to be tackled. The first challenge is finding a good quality data set to train the ML model, as a model can only be as good as its training data.[chen:AdvEnergyMat2020, raccuglia:Nature2016, ma:NPJCompMat2024] To tackle this challenge, a consistent and qualitative data set is built in-house. The second challenge is overcoming the data hungry nature of typical ML methods. Since the goal is to make synthesis optimization more efficient using ML, it makes no sense to first conduct hundreds of experiments that can serve as data to train the model. Therefore, a small data set of only a few experiments is used.[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022] Finally, within the context of lab-scale experiments, we emphasize interpretable ML: the model should not only predict outcomes, but also provide insight into the role of different synthesis parameters.
@@ 2 Methodology
Within this work, we follow a four step workflow (Figure 1), of which the latter three can form a closed-loop experimentation methodology. The initial set of parameters for the data set is selected using Latin Hypercube Sampling (LHS) to achieve optimal parameter space coverage.[borisut:Processes2023] The sampled syntheses are performed in the lab via MW-assisted polyol synthesis. The resulting dispersions are analyzed via dynamic light scattering (DLS) to determine the particle size of the Cu NPs, while the surface plasmon resonance (SPR) peak is determined using UV-Vis measurements. The effective synthesis parameters111These are the true experimental values corresponding to the LHS. and corresponding particle sizes are combined into the data set for the ML modeling. In a closed-loop approach, the results of the ML models can be used to assist in further experimental syntheses, or further validation of the ML model.
@@ 2.1.2 Sampling synthesis parameters
LHS is used to obtain an optimized data set with minimal correlation between the various experiments while providing an optimal coverage of the parameter space under investigation.[teichert:DaltonTrans2018, borisut:Processes2023, braham:DaltonTrans2020] In this work, the parameter space comprises three dimensions:
@@
Using LHS, 25 experiments are selected that fill this parameter space as optimally as possible.
@@ 2.1.3 Cu NPs synthesis
The Cu NPs are synthesized via the microwave (MW)-assisted polyol route[fievet:ChemSocRev2018, blosi:JournalNPResearch2010, teichert:DaltonTrans2018, carroll:JournalPhysChemC2011, torras:CrystalGrowth2021] in a CEM Discover SP MW reactor. First, the amount of is weighed into the reaction vial. It is then dissolved in 5 ml of ethylene glycol under sonication. After the Cu precursor has completely dissolved, the tetramethylammonium hydroxide (TMAH) is added to the solution and further mixed and degassed under sonication. TMAH is consistently used at a molar ratio of relative to the Cu precursor. The reaction vessel is then placed in the microwave reactor, with the head space filled with inert argon up to a pressure of 5 bar, and is first stirred for minutes at to ensure complete dissolution of the precursor. Afterwards, the temperature is raised to the desired reaction temperature with continuous stirring and maintained for the desired reaction time. After the reaction time has passed, the reaction is quenched by actively cooling the vessel with compressed air. A brown-red suspension is obtained, which is typical for Cu NPs.[blosi:JournalNPResearch2010] To collect the Cu NPs, the reaction mixture is diluted with ml of ethanol and centrifuged at rpm for minutes. Subsequently, the supernatant is removed, and the resulting NPs are again dispersed in ml of ethanol.
@@ 2.2 Computational modeling
Two modeling tasks are performed on the data set: a regression task based on ensemble regression ML, implemented in the AMADEUS framework,[AMADEUS2020, vanpoucke:JournalApplPhys2020] and these results are compared with classical DoE[shahriari:ProceedingsIEEE2015], to gain insight in the potential of ML.
@@ 2.2.1 Ensemble regression models
Since ML using small data sets is very sensitive to the selected training data, and hyperparameter tuning suffers from this volatility[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022], we make use of the scikit-learn-based AMADEUS framework which was designed to deal with small data sets (Figure 2).[vanpoucke:JournalApplPhys2020, vanpoucke:PolymerInt2022, pedregosafabianJournalMLResearch2011, AMADEUS2020] Pasting-type ensemble models consisting of base instances are created, using either linear or polynomial base models in combination with LASSO regularization.[tibshirani:RoyalStatSocB1996] Each of the instances is trained on a different randomly selected 80/20 split (in-bag/out-of-bag (OOB)) of the complete data set. The resulting final ensemble model is a single base instance equal to the average of the 100 base model instances.
@@
Data set A: All data points are included in the data set, and the predicted particle size is the smallest particle size present in the DLS particle-size distribution.
@@
Data set B: All data points are included in the data set, and the predicted particle size corresponds to the peak with the highest percentage contribution in the DLS particle-size distribution.
@@
Data set C: Only the data points with a mono-modal DLS particle-size distribution are included in the data set.
@@
A further measure employed to reduce overfitting, is consideration of what one could call ensemble importance.[vanpoucke:PolymerInt2022, VerdingP:AdvMaterTech2025] The AMADEUS framework tracks how often a polynomial feature is retained (i.e. has non-zero coefficient) within the ensemble, after regularization of the features in each of the model instances. This provides a measure of importance of the features with regard to a well-generalizing model. A feature with an ensemble importance of % has a non-zero coefficient in all model instances in the ensemble. If a feature has an ensemble importance of %, this means that in % of the model instances in the ensemble, the coefficient for this feature is zero. Based on the trends in ensemble importance, seven features are selected (cf. SI Figure 6):
@@ 2.2.2 Classical statistical methods (DoE)
Using DoE[shahriari:ProceedingsIEEE2015] software JMP Pro (SAS Institute Inc.), a linear regression model is fitted for data set C, including terms up to the second degree. Insignificant terms are removed until only significant terms remain. In contrast to machine learning approaches, the available data set is not split into training and test data; all data points are used to fit the regression model. The performance of the model is determined by calculating the expected size for all data points using the equation obtained, and thus should be compared to quality estimates on the in-bag data set.
@@ 3.1 Cu NP syntheses
The DLS measurements show a peak above nm for five samples. However, these peaks are not included in the analysis. Due to their very low intensity (), it can be assumed that these results are due to contamination, such as dust particles, during the measurements. The synthesis yields Cu NPs with a particle size between nm and nm. The DLS shows a bi-modal and tri-modal particle-size distribution for five and two samples, respectively. The remaining samples show a mono-modal particle-size distribution. Analysis of the DLS intensity-size plots allows us to identify the smallest particle size present. The median of these particle sizes is nm over all samples.
Based on the SPR location, the effective formation of Cu metal () NPs is verified for the different samples. The different UV-Vis spectra show a clear SPR peak between nm and nm for most samples, which is a strong indication of the presence of NPs.[blosi:JournalNPResearch2010] Samples , , and do not show a clear SPR feature, which may be due to low particle concentration, oxidation[vargas-urbano:Materials2022], or the presence of ultra small particles (), which do not exhibit a well-defined SPR; for the other samples, at least some is present. Given the limited yield of the syntheses, no X-ray diffraction measurements could be performed to provide a conclusive answer, but the combination of the red-brown color of the solution and the presence of a clear SPR peak in the UV-Vis spectra strongly indicate that NPs have indeed formed. To validate this conclusion, Selected Area Electron Diffraction analysis is conducted on sample 22 using a transmission electron microscope operated at an accelerating voltage of 120 kV. The resulting diffraction rings are indexed to the face-centered cubic structure of copper, with prominent rings corresponding to the (111), (200), and (220) planes. The primary diffraction rings are assigned the following spacings: Å for the (111) plane, Å for the (200) plane, and Å for the (220) plane. These values agree with the standard values for FCC copper (JCPDS card no. 04-0836)[icdd2001], confirming the formation of NPs.

### v2

@@ Abstract
Copper nanoparticles (Cu NPs) have a broad applicability, yet their synthesis is sensitive to subtle changes in reaction parameters. This sensitivity, combined with the time- and resource-intensive nature of experimental optimization, poses a major challenge in achieving reproducible and size-controlled synthesis. While Machine Learning (ML) shows promise in materials research, its application is often limited by scarcity of large high-quality experimental data sets. This study explores ML to predict the size of Cu NPs from microwave-assisted polyol synthesis using a small data set of 25 in-house performed syntheses. Latin Hypercube Sampling is used to efficiently cover the parameter space while creating the experimental data set. Ensemble regression models successfully predict particle sizes with high accuracy (), outperforming classical statistical approaches (). Additionally, classification models using both random forests and Large Language Models (LLMs) are evaluated to distinguish between large and small particles. While random forests show moderate performance, LLMs offer no significant advantages under data-scarce conditions. Overall, this study demonstrates that carefully curated small data sets, paired with robust classical ML, can effectively predict the synthesis of Cu NPs and highlights that for lab-scale studies, complex models like LLMs may offer limited benefit over simpler techniques.
@@ 1 Introduction
Metal nanoparticles (NPs) play a central role in a broad range of applications, including catalysis, electronics, and biomedicine, due to their unique size- and shape-dependent properties.[14, 48, 18, 30] However, controlling these properties requires very precise control of various NP characteristics, including their morphology (i.e., size and shape), composition, and surface chemistry down to the nanometer scale. Achieving this level of control remains one of the biggest challenges in NP synthesis, as small variations in the synthesis can lead to final particles with entirely different properties.[18, 4] Traditionally, synthesis optimization relies on iterative, empirical approaches that are slow, resource-intensive, and often lack mechanistic insight.[4] To overcome these limitations, more systematic, data-driven strategies are an interesting route to explore. Indeed, recent advances in artificial intelligence (AI) and machine learning (ML) offer promising opportunities to accelerate the development of reproducible and tailored NP syntheses.[4, 5, 9, 35, 34, 21, 2]
@@
By uncovering hidden correlations in complex parameter spaces, ML can guide experimental design and reduce the number of required experiments.[9, 35, 34, 21] In materials science, ML is now routinely used to uncover the structure–property relationships of materials, thereby accelerating the discovery of new materials with specific properties.[9, 35, 34, 21] However, for synthesis optimization, the key challenge lies in predicting how synthesis parameters determine the resulting crystal structure or particle morphology. These synthesis–structure relationships are far less explored, yet crucial for rational synthesis design. As a result, recent data-driven approaches are shifting towards models that learn synthesis–structure relationships to accelerate synthesis optimization.[2, 42, 47] Unlike traditional one-variable-at-a-time (OVAT) approaches, which are slow and fail to capture the multidimensional correlations that govern synthesis outcomes, ML can account for such interactions and thereby accelerate synthesis optimization.[47, 26] However, while large data sets underpin many ML successes [23, 28, 13], they are rarely feasible in laboratory-scale chemistry, where the creation of each data point requires a significant amount of time and resources. In the specific case of inorganic NP synthesis, literature-based data sets are often inconsistent and biased, as failed experiments remain unpublished[11, 32, 27], as shown by Huo et al.[23], who expected up to % of the values to be incorrect. As such, robust ML workflows for small, high-quality data sets often remain underexplored.
@@
ML can identify which synthesis parameters most strongly influence inorganic NP outcomes, as demonstrated by Guda et al.[19], who trained ML models on 30 data points to classify the results of gold NP synthesis (success, shape, and formation kinetics) based on the precursor concentrations, reducing agent, and surfactant. ML also enables closed-loop experimentation, where ML models suggest subsequent experiments based on previous results, accelerating optimization and supporting high-throughput synthesis.[37, 15] Pellegrino et al.[31] trained neural networks (NNs), with 20 initial data points, to predict the size, dispersion, and aspect ratio of NP, identifying optimal synthesis parameters through reverse engineering. Similarly, Williams et al.[46] combined Design of Experiments (DoE) and random forests to accurately predict the yield of Ag nanowires in a polyol-based flow reactor, based on 31 samples.
@@
In the current work, the prediction of copper (Cu) NP size as function of synthesis parameters is explored. Cu NPs hold significant promise for various catalytic processes, such as electro-catalysis and photo-catalysis, due to their abundance, cost-effectiveness, and versatile properties.[20, 33, 18] For example, Reske et al.[33] showed that decreasing the size of Cu NPs significantly improves their catalytic activity. Different synthesis methods give rise to different types of Cu NPs, which may differ in particle size, morphology, and extent of oxidation. While metallic Cu NPs are nominally , they often undergo (partial) surface oxidation depending on the synthesis conditions, leading to phases of varying composition.[18] The wide variety of possible synthesis outcomes, as well as the large parameter space, complicates the development of robust synthesis procedures to obtain specific properties.
@@
Though ML is a tool with great potential for Cu NP synthesis, two challenges need to be tackled. The first challenge is finding a good quality data set to train the ML model.[11, 32, 27] To tackle this challenge, a consistent and qualitative data set is built in-house. The second challenge is overcoming the data hungry nature of typical ML methods. Since the goal is to make synthesis optimization more efficient using ML, relying on hundreds of preliminary experiments to generate training data would be impractical in a lab-scale setting, both in terms of time and resource consumption. Instead, we focus on constructing a carefully selected, information-rich dataset that maximizes experimental variation while minimizing experimental effort.[29, 42] Finally, within the context of lab-scale experiments, we emphasize interpretable ML: the model should not only predict outcomes, but ideally also provide insight into the role of different synthesis parameters.
@@
Recent work by Zheng et al.[49], which applies large language models (LLMs) to synthesis protocols reported in the MOF literature, served as inspiration to examine the potential of LLMs for synthesis prediction. Motivated by this, we investigate whether LLMs can provide meaningful predictive performance in a small-data experimental setting and benchmark their performance against a classic ML approach.
@@ 2 Methodology
Within this work, we follow a four step workflow (Figure 1), of which the latter three can form a closed-loop experimentation methodology. The initial set of parameters for the data set is selected using Latin Hypercube Sampling (LHS) to achieve optimal parameter space coverage.[7] The sampled syntheses are performed in the lab via MW-assisted polyol synthesis. The resulting dispersions are analyzed via dynamic light scattering (DLS) to determine the particle size of the Cu NPs, while the surface plasmon resonance (SPR) peak is determined using UV-Vis measurements. The effective synthesis parameters[1] and corresponding particle sizes are combined into the data set for the ML modeling. In a closed-loop approach, the results of the ML models can be used to assist in further experimental syntheses, or further validation of the ML model.
@@ 2.1.2 Sampling synthesis parameters
LHS is used to obtain an optimized data set with minimal correlation between the various experiments while providing an optimal coverage of the parameter space under investigation.[38, 7, 8] In this work, the parameter space comprises three dimensions:
@@
Using LHS, 25 experiments are selected that fill this parameter space as optimally as possible. (cf. SI Table V)
@@ 2.1.3 Cu NPs synthesis
The Cu NPs are synthesized via the microwave (MW)-assisted polyol route[17, 6, 38, 10, 40] in a CEM Discover SP MW reactor. First, the amount of is weighed into the reaction vial. It is then dissolved in 5 ml of ethylene glycol under sonication. After the Cu precursor has completely dissolved, the tetramethylammonium hydroxide (TMAH) is added to the solution and further mixed and degassed under sonication. TMAH is consistently used at a molar ratio of relative to the Cu precursor. The reaction vessel is then placed in the microwave reactor, with the head space filled with inert argon up to a pressure of 5 bar, and is first stirred for minutes at C to ensure complete dissolution of the precursor. Afterwards, the temperature is raised to the desired reaction temperature with continuous stirring and maintained for the desired reaction time. After the reaction time has passed, the reaction is quenched by actively cooling the vessel with compressed air. A brown-red suspension is obtained, which is typical for Cu NPs.[6] To collect the Cu NPs, the reaction mixture is diluted with ml of ethanol and centrifuged at rpm for minutes. Subsequently, the supernatant is removed, and the resulting NPs are again dispersed in ml of ethanol.
@@ 2.2 Computational modeling
Two modeling tasks are performed on the data set: a regression task based on ensemble regression ML, implemented in the AMADEUS framework,[43, 29] and these results are compared with classical DoE[36], to gain insight in the potential of ML.
@@ 2.2.1 Ensemble regression models
Since ML using small data sets is very sensitive to the selected training data, and hyperparameter tuning suffers from this volatility[29, 42], we make use of the scikit-learn-based AMADEUS framework which was designed to deal with small data sets (Figure 2).[29, 42, 16, 43] Pasting-type ensemble models consisting of base instances are created, using either linear or polynomial base models in combination with LASSO regularization.[39] Each of the instances is trained on a different randomly selected 80/20 split (in-bag/out-of-bag (OOB)) of the complete data set. The resulting final ensemble model is a single base instance equal to the average of the 100 base model instances.
@@
Data set A: All data points are included in the data set, and the predicted particle size is the smallest particle size present in the DLS particle-size distribution. (cf. SI Table VI)
@@
Data set B: All data points are included in the data set, and the predicted particle size corresponds to the peak with the highest percentage contribution in the DLS particle-size distribution. (cf. SI Table VII)
@@
Data set C: Only the data points with a mono-modal DLS particle-size distribution are included in the data set. (cf. SI Table VIII)
@@
A further measure employed to reduce overfitting, is consideration of what one could call ensemble importance.[42, 45] The AMADEUS framework tracks how often a polynomial feature is retained (i.e. has non-zero coefficient) within the ensemble, after regularization of the features in each of the model instances. This provides a measure of importance of the features with regard to a well-generalizing model. A feature with an ensemble importance of % has a non-zero coefficient in all model instances in the ensemble. If a feature has an ensemble importance of %, this means that in % of the model instances in the ensemble, the coefficient for this feature is zero. Based on the trends in ensemble importance, seven features are selected (cf. SI Figure 6):
@@ 2.2.2 Classical statistical methods (DoE)
Using DoE[36] software JMP Pro (SAS Institute Inc.), a linear regression model is fitted for data set C, including terms up to the second degree. Insignificant terms are removed until only significant terms remain. In contrast to machine learning approaches, the available data set is not split into training and test data; all data points are used to fit the regression model. The performance of the model is determined by calculating the expected size for all data points using the equation obtained, and thus should be compared to quality estimates on the in-bag data set.
2.2.3 Classification analysis
A binary classification is performed on the particle size using data set A. (cf. SI Table VI) The aim is to predict whether the resulting particles will be large or small based on the synthesis parameters. Particles are defined as large if they are larger than the median size in the data set; if they are smaller than the median size, they are defined as small. For this classification task, the performance of complex LLMs is compared with that of a simple random forest classifier using accuracy and kappa () as performance metrics. indicates the agreement between predictions and actual labels, but corrects for coincidental agreements. It therefore indicates how well the model performs compared to random guessing. If there is a perfect match between predictions and actual labels, while indicates a performance equal to random guesses. If the model performs even worse than random guessing.
Two different LLMs are trained and compared with each other for the binary classification task: GPT-J and LLAMA 3.1.[25, 41], to predict the outcome as large particles (‘1’) or small particles (‘0’) based on the experimental input parameters. The LLMs are fine-tuned using the following prompt as input:
‘What is the size of nanoparticles synthesized with a mol/litre Copper Acetate precursor concentration, a mol/liter tetramethylammonium hydroxide concentration, heated for minutes at degrees celcius?’
In this prompt, acts as a placeholder for the reaction parameter value. The LLM predicts ‘0’ or ‘1’ corresponding to experimentally small or large NPs, respectively. The training data set is limited to a maximum of examples, while the test data set remained constant across all runs (e.g., examples). The number of epochs[12] is first evaluated at , , , and . For each combination of training size and epoch count, five independent runs are performed to compute average performance metrics. GPT-J is used for this screening. Because the task involves binary classification of a balanced data set, any accuracy above % indicates performance better than random guessing.
In comparison to the LLMs, an ensemble approach (e.g., random forest models) is used for binary classification. Automatic hyperparameter selection is used to determine the optimal value for the number of trees within a forest (), and the maximum depth of a tree (). Based on this hyperparameter tuning, is set to and to 3. With these hyperparameters, a single random forest is trained on a random subset of the data (80/20 train/test split). Repeating this process ten times, shows a high dependency on which specific data points are include in the training (and test) data, in line with previous observations with the volatility of results for small data ML.[29] Our results show the models accuracy varying between and . Because of this, an ensemble of random forests is trained, each random forest on a new random subset of the data. Each generated model is tested on its OOB data, and these predictions are logged. After training and evaluating all models, the OOB predictions of all models are used to obtain a final classification through hard voting. For the OOB predictions, the number of times they are classified as small (‘0’) and the number of times as large (‘1’) are counted for each data point. The class with the majority of votes will be the final predicted class.
@@ 3.1 Cu NP syntheses
The DLS measurements show a peak above nm for five samples. However, these peaks are not included in the analysis. Due to their very low intensity (), it can be assumed that these results are due to contamination, such as dust particles, during the measurements. The synthesis yields Cu NPs with a particle size between nm and nm. The DLS shows a bi-modal and tri-modal particle-size distribution for five and two samples, respectively. The remaining samples show a mono-modal particle-size distribution. (cf. SI Figures 9, 10 and 11) Analysis of the DLS intensity-size plots allows us to identify the smallest particle size present. The median of these particle sizes is nm over all samples. (cf. SI Figure 7)
Based on the SPR location, the effective formation of Cu metal () NPs is verified for the different samples. The different UV-Vis spectra show a clear SPR peak between nm and nm for most samples, which is a strong indication of the presence of NPs.[6] Samples , , and do not show a clear SPR feature, which may be due to low particle concentration, oxidation[44], or the presence of ultra small particles (), which do not exhibit a well-defined SPR; for the other samples, at least some is present. Given the limited yield of the syntheses, no X-ray diffraction measurements could be performed to provide a conclusive answer, but the combination of the red-brown color of the solution and the presence of a clear SPR peak in the UV-Vis spectra (cf. SI Figures 13 and 14 and SI Table IX) strongly indicate that NPs have indeed formed. To validate this conclusion, Selected Area Electron Diffraction analysis is conducted on sample 22 using a transmission electron microscope operated at an accelerating voltage of 120 kV. The resulting diffraction rings are indexed to the face-centered cubic structure of copper, with prominent rings corresponding to the (111), (200), and (220) planes. (cf. SI Figure 8) The primary diffraction rings are assigned the following spacings: Å for the (111) plane, Å for the (200) plane, and Å for the (220) plane. These values agree with the standard values for FCC copper (JCPDS card no. 04-0836)[24], confirming the formation of NPs.